# Assessing the accuracy of the star formation rate measurements by direct star count in molecular clouds

### v2

@@ 2.3 The star formation history
When deriving the SFR using Eq. 1, and in almost all observational studies, the assumption is made that the SFH in a star-forming region has been constant over the entire lifetime of the regions (Evans et al. 2009 and references therein). This assumption may be true in some molecular clouds, possibly in low-mass clouds where star formation proceeds at a slow pace, but is unlikely to be true for all clouds. The age distribution of stars in young clusters is not always a flat distribution as exemplified by the case of the Orion Nebula Cluster (ONC, Dib et al. 2013). In the ONC, the age distribution of stars is well reproduced by the sum of a linearly increasing SFH at old ages (up to 6 Myr) and an acceleration (i.e., a burst) of star formation at younger ages (see Figure 13 in Dib et al. 2013). Similar SFHs consisting of a slow increase when clouds are in the process of assembling followed by a burst of star formation at later times as most of the gas sinks to the central regions of the clouds, are also observed in numerical simulations (e.g., Vázquez-Semadeni et al. 2017, Guszejnov et al. 2022). It has also been shown that such burst-like SFHs better describe the SFH of clusters such as the Orion Nebular cluster (Kounkel 2020; see also Megeath et al. 2022), and are necessary in order to reproduce the Mass-Luminosity relation of massive star forming regions (Zhou et al. 2024c,d). In this work, we test both assumptions. We consider cases where the SFH is constant and others where it varies over time. For cases with a constant SFH (labelled as constant or flat), we consider a fiducial case of with a birthrate of stars yr-1 (i.e., 500 stars form over a period of yr). For the time-dependent SFH, a variety of functional forms can be used to describe a slowly increasing SFH followed by a burst. For simplicity, we use a Gaussian function, for which we are able to vary the position of the peak and the width of the distribution. The SFH is described by a probability for the time of birth of the stars, , and in the case of a time-varying SFH, this is given by:
@@ 3.1 The effect of the SFH
First, we present results from two fiducial cases, one with a constant SFH and another one with a Gaussian-like SFH. For these models, we consider a cluster who will end up containing stars and for which the stars are sampled from a Milky-Way like single-star IMF (i.e., , , and M⊙). For each case we generate 250 models. Figure 1 displays four realizations of the Milky Way-like randomly sampled IMF. Figure 2 and Fig. 3 each display four realizations of the birth times of protostars. For the case of the flat SFH, displayed in Fig. 2, the birth time of protostars are randomly selected with a uniform probability in the time span between and , whereas for the cases of the Gaussian-like SFH displayed in Fig. 3, the birth times of protostars are randomly sampled using Eq. 4. The next step is to assign to each protostar lifetimes that are associated with each protostellar class. As described above, these lifetimes are randomly drawn using the probability distribution functions described by Eq. 3. As fiducial values, we adopt here those listed in §. 2.2 namely Myr and Myr for Class 0, Myr and Myr for Class I, Myr and Myr for the F Class, Myr and Myr for Class II, and Myr and Myr for Class III. As an example, Fig. 3 displays the distribution functions of the protostellar lifetimes for all YSOs that form in one cluster. When randomly sampling each YSO lifetime from the Gaussian function shown in Eq. 3, the randomly sampled values are drawn in the range of the uncertainty of each quantity.
@@
With the above input quantities, we track the number of YSOs found in each class as a function of time. Figure 4 (left panel) displays four examples (out of 250 generated models) for the case with a flat SFH, a Milky-like IMF, and in each cluster. Figure 4 indicates that within the temporal fluctuations due to the random sampling, the total number of Class 0 protostars is constant over time. The number of Class I and Class F protostars starts to increase after and Myr, respectively, and both reach a plateau after Myr. The numbers of Class II and Class III PMS stars display the same behavior but at later times (i.e., after 2 to several Myr). The right panel in Fig. 4 displays the time evolution of the number of YSOs in each Class for four cases with a Gaussian-like SFH (with the fiducial parameters). This figure clearly shows the peak in Class 0 protostars at Myr and this peak shifts to later epochs for more evolved protostars. Note that in these cases, the number of Class II and Class III PMS stars also declines at later times because most of the protostars where formed early at Myr and almost no new protostars were formed after Myr.
@@
Using the number of YSOs found in each Class and at each epoch as well as the mean values of their lifetimes (i.e., those suggested by Evans et al. 2009), we can now apply Eq. 1 and calculate the corresponding SFRs. We also measure the true SFR in each model. For this, we take a timestep of yr. This guaranties that the smallest protostellar lifetimes are well resolved. Then we simply calculated the true SFR as being the total mass of protostars that form within each time step
@@
where runs over all stars that form between and . Fig. 5 (left) displays the time evolution of the SFR for four cases where the SFH is flat. We recall that this is usually the assumption that is made in the observations when deriving the SFR. Figure 6 (left panel) displays the true SFR measured directly from the models (dashed lines) and an SFR estimate using the total number of YSOs (full line). For the latter, a characteristic YSO lifetime of 2.5 Myr was used (Pokhrel et al. 2020). What the left panels of Fig. 5 and Fig. 6 show is that while the SFR estimates measured using a single Class converge to the real value, they do so with a delay that increases for more evolved YSOs. This implies that for cases where the SFH is constant, measurements of the SFR will only be reliable if only Class 0 and I protostars are considered in young star forming regions and Class II PMS stars in older regions (i.e., ages Myr). Figure 6 (left panel) additionally shows that using the total number of YSOs provides a very poor approximation to the real SFR. It underestimates the true SFR by a factor of to a in young star-forming regions and overestimates the true SFR by a factor of a few to several for older regions. The right panels in Fig. 5 and Fig. 6 display the same SFRs calculated for cases where the SFH is a Gaussian-like function with the fiducial values of the parameters, Myr and Myr. The same patterns that are observed with a constant SFH are also prevalent here, with the additional effect that at advanced epochs (i.e., ages Myr), the estimates of the SFR that are based either on the Class II and Class III PMS stars or the ones based on the total number of YSOs overestimate the true SFRs by several orders of magnitudes. This is due to the fact that while star formation has ceased, Class II and Class III PMS stars are long lived and their numbers bias the measurement of the SFR.
@@
In order to get a clearer picture of the offset between the SFRs estimated using the YSO populations and the true SFR, we show in Fig. 7 the ratio between these quantities. In both panels of this figure, the ratios were obtained by averaging over the 250 realizations (i.e., 250 clusters) each with and a Milky Way-like IMF. The top panel correspond to cases with a flat SFH and the bottom panel to cases with the Gaussian-like SFH with all clusters in the latter case having Myr and Myr. The light dashed line in both panels corresponds to the 1 Poisson uncertainty on each measurements. Ideally, for these measurements to be accurate, the offset parameter =SFRi/SFRreal would have to be close to the unity. However, this is far from being the case with the worst departure from unity corresponding to the time-dependent, Gaussian-like SFH, especially around and after the peak of star formation at Myr. For these time-dependent models of the SFH, we expect the level of departure from unity to depend of the duration of the burst (i.e., the width of the Gaussian). We therefore run additional models, each with 250 clusters, of Gaussian-like SFHs. Since our fiducial case has a value of Myr, we consider the additional cases where Myr (a narrow, high-amplitude burst) and Myr (an extended, low amplitude burst). The results for these models are displayed in Fig. 8 for the narrow burst (top panel) and extended burst (bottom panel). Qualitatively, both cases display the same features and are similar to the fiducial case displayed in Fig. 7 (top panel). However, a noticeable difference for the narrow burst case (i.e., case with Myr) is that all SFR measurements, with the exception of the one based on the Class 0 population, using either the other Classes of YSOs or their total number fail to reproduce the true SFR.
@@
The conclusion that can be drawn from Fig. 7 and Fig. 8 is that the only reliable estimate of the SFR at all epochs is the one based on the population of Class 0 protostars. The latter measurement is always accurate, irrespective of the SFH. Measurements of the SFR based on the population of more evolved YSOs can provide reliable estimates of the SFR in cases where the SFH is constant or displays a weak burst, and this is only valid when the star-forming regions are themselves more evolved (i.e., ages of 2 Myr or older, and harbor a significant fraction of evolved YSOs). In all circumstances, estimates of the SFR that rely on the total number of YSOs and a characteristic YSO lifetime (in this work Myr) are not accurate and this is true irrespective of the SFH and the considered age of the region. Finally, we should mention that we have performed models with lower and higher numbers of YSOs, namely and , respectively while keeping Myr in all models. Models with lower or higher number of protostars display the same effects as those observed in models with YSOs but with an increased and decreased level of temporal fluctuations, respectively (see App. A).
@@ 3.2 The effect of the protostellar mass function
Recent studies suggest a non-negligible degree of variation in the parameters that characterize the shape of the IMF of young clusters in the Milky Way and in M31 (Dib 2014; Weisz et al. 2015; Lim et al. 2015; Dib et al. 2017, Dib 2023b). Dib et al. (2017) found that an intrinsic scatter in the IMF parameters is necessary in order to match the observed fraction of single O stars in Galactic young clusters (cluster ages Myr). For a TPL function representing the system IMF, Dib et al. (2017) found that the standard deviation in , , and is , , and M⊙. Here, as we are dealing with a single star protostellar mass function, we model the distribution of each of the mass function parameters with a Gaussian centered around the Galactic values (, , and M⊙) and with a standard deviation on each parameter that is of its mean value, namely , , and M⊙. Similar to the previous calculations, we generate a sample of 250 clusters each with and where the parameters of the IMF each cluster are randomly drawn from these Gaussian distributions. The IMF parameters are randomly drawn from their Gaussian distribution each in the uncertainty range. Figure 9 displays the cases for four clusters with such a variable IMF. We then follow the same procedure as above and derive the SFRs based on the different YSO classes as well as the one based on the total YSO population. Adopting the case of a flat SFH, Fig. 10 displays the time evolution of these SFRs normalized by the true SFR. Comparing this figure with Fig. 7 (top panel), we conclude that a variable IMF has little effect on the derived SFRs. The same features observed in the top panel of Fig. 7 and where the IMF of clusters are randomly drawn from a Milky Way-like IMF are seen in Fig. 10, namely a faithful reproduction of the true SFR (SFRreal) by the SFR based on the Class 0 population, and a reproduction of the SFRreal by the SFRs based on the more evolved populations (Class I and beyond) at more advanced epochs. The SFR derived using the total YSO population remains a poor approximation to the true SFR as it severely underestimates SFRreal at early epochs and overestimates it at later epochs.
@@ 4 Constraining the SFH using protostellar counts and SFR estimates
In §. 3, we argued that the SFR derived using the Class 0 YSO count provides a reliable estimate of the true SFR in star-forming regions, irrespective of the SFH. However, as exemplified by the models with stars, the number of Class 0 YSOs remains quite modest at most if not all epochs. Furthermore, Class 0 YSOs are sometimes difficult to detect as they are deeply embedded in the clouds (Evans et al. 2009). In contrast, the numbers of evolved YSOs is much larger. The number of Class II YSOs is a factor times larger than that of Class 0 YSOs. For this reason, using Class II object to derive the SFR can be an appealing option in order to make use of the Class II larger number of objects. The caveat to using this argument is that the offset between the Class II derived SFR (SFRII) and the true SFR depends on the SFH. As shown in §. 3, for the case of a flat SFH, the offset between SFRII and SFRreal exists only at early epochs ( 2 Myr from the start of star formation). For the case of burst-like SFH, the situation is more complex, and an offset between SFRII and SFRreal exists at all times. It is possible to reverse the arguments and assumes that the observed SFRII corresponds to the SFRreal and verify whether this assumption hold for the predictions of constant SFH or a burst-like one. Figure 11 displays the Class I and Class II SFR offsets, plotted against the ratio and , respectively, for the case with a flat SFH. Figure 12 is similar to Fig. 11 but for the fiducial case with a Gaussian-like SFH. The dependencies between the offsets and and the number ratios are noticeably different between these cases with a different SFH.
@@
We can test the usefulness of these relations in discriminating between SFHs by including observational data. We use the data presented in Table 3 and 5 of Evans et al. (2009) for five nearby molecular clouds, namely, Cha II, Lupus, Perseus, Serpens, and Ophiucus. Table 5 lists the numbers of YSOs in Class I, the F Class, Class II, and Class III. Since the numbers of Class II YSOs are larger than those in all other classes, this indicates that star formation have started at least Myr ago which is roughly equal plus a fraction of the Class II lifetime. Table 3 of Evans et al. (2009) also lists SFR estimates for these clouds. These SFR measurements are derived using the Class II YSO populations and were presented in separate works but are all based on c2d observations. Using the number counts of Class I and Class II YSOs, we use Eq. 1 to calculate SFRI and SFRII and consequently and . If the assumption that the SFR values quoted for this clouds are the true values, the ratio for these clouds should coincide with the prediction of the model with a constant SFR. Looking at Fig. 11 (bottom panel) confirms that this is indeed the case. Furthermore, the upper panel of Fig. 11 shows that the vs. relation for these five clouds is also reasonably well matched by the model with a constant SFH. In contrast, the model with a burst-like SFH does not overlap with the observational data as can be seen in the lower and upper panels of Fig. 12. The conclusion that can be drawn from Fig. 11 and Fig. 12 is that the SFH in these five, nearby, and low to intermediate mass clouds (from to M⊙; Table 1 in Evans et al. 2009 and references therein) is relatively constant. It would be interesting to include data points for more massive star-forming regions and regions that are located further away in the Galaxy.
@@
and for the time-dependent, Gaussian-like SFH:
As pointed out in §. 2.2 the Class III YSOs observed by Spitzer for these nearby star forming regions are incomplete and Eqs. 6 and 7 should be revised with more complete class III censuses. More complete censuses of Class III YSOs could be provided by X-ray observations as shown by Preibisch et al. (2005) in the Orion Nebular Cluster. However, we note that a more accurate accounting for Class III YSOs will reinforce our conclusions concerning the comparison of our models with the observations. Accounting for a higher number of missing Class III YSOs would lead the observational number ratios and in Figs. 11 and 12 to move to the left. This would not affect the agreement between the flat SFH models and the observations but it would worsen the case of burst-like models compared to the observations. While it is not yet possible to include a more massive region such as the entire Orion cloud due to the fact that the census of Class II YSOs is incomplete (Furlan et al. (2016)444Megeath et al. (2012,2016) performed a comprehensive survey of the YSO population in the Orion clouds. However the YSOs in these works were categorized in the broad category of protostars and PMS stars and for this reason, the Orion data points could not be includedin Figs. 11 and 12, these authors found that the number of Class 0 YSOs in Orion (). Using Eq. 1, we conclude that the true SFR for the Orion of M⊙ yr-1. This value sits in the middle between the SFRs measured by Lada et al. (2010) for the Orion A and Orion B clouds.

### v4

@@ 2.3 The star formation history
When deriving the SFR using Eq. 1, and in almost all observational studies, the assumption is made that the SFH in a star-forming region has been constant over the entire lifetime of the regions (Evans et al. 2009 and references therein). This assumption may be true in some molecular clouds, possibly in low-mass clouds where star formation proceeds at a slow pace, but is unlikely to be true for all clouds. The age distribution of stars in young clusters is not always a flat distribution as exemplified by the case of the Orion Nebula Cluster (ONC, Dib et al. 2013). In the ONC, the age distribution of stars is well reproduced by the sum of a linearly increasing SFH at old ages (up to 6 Myr) and an acceleration (i.e., a burst) of star formation at younger ages (see Figure 13 in Dib et al. 2013). Similar SFHs, consisting of a slow increase when clouds are in the process of assembling followed by a burst of star formation at later times as most of the gas sinks to the central regions of the clouds, are also observed in numerical simulations (e.g., Vázquez-Semadeni et al. 2017, Guszejnov et al. 2022). It has also been shown that such burst-like SFHs better describe the SFH of clusters such as the Orion Nebular cluster (Kounkel 2020; see also Megeath et al. 2022), and are necessary in order to reproduce the mass–luminosity relation of massive star-forming regions (Zhou et al. 2024c,d). In the present work, we test both assumptions. We consider cases where the SFH is constant and others where it varies over time. For cases with a constant SFH (labelled as constant or flat), we consider a fiducial case with a birthrate of stars yr-1 (i.e., 500 stars form over a period of yr). For the time-dependent SFH, a variety of functional forms can be used to describe a slowly increasing SFH followed by a burst. For simplicity, we use a Gaussian function, for which we are able to vary the position of the peak and the width of the distribution. The SFH is described by a probability for the time of birth of the stars, , and in the case of a time-varying SFH this is given by
@@ 3.1 The effect of the SFH
First, we present results from two fiducial cases, one with a constant SFH and another with a Gaussian-like SFH. For these models, we consider a cluster that will end up containing stars sampled from a Milky Way-like single-star IMF (i.e., , , and M⊙). For each case, we generate 250 models. Figure 1 displays four realizations of the Milky Way-like randomly sampled IMF. Figure 2 and Fig. 3 each display four realizations of the birth times of protostars. For the case of the flat SFH displayed in Fig. 2, the birth times of protostars are randomly selected with a uniform probability in the time span between and , whereas for the cases of the Gaussian-like SFH displayed in Fig. 3, the birth times of protostars are randomly sampled using Eq. 4. The next step is to assign to each protostar lifetimes that are associated with each protostellar class. As described above, these lifetimes are randomly drawn using the probability distribution functions described by Eq. 3. As fiducial values, we adopt here those listed in §. 2.2, namely Myr and Myr for Class 0, Myr and Myr for Class I, Myr and Myr for the F Class, Myr and Myr for Class II, and Myr and Myr for Class III. As an example, Fig. 3 displays the distribution functions of the protostellar lifetimes for all YSOs that form in one cluster. When randomly sampling each YSO lifetime from the Gaussian function shown in Eq. 3, the randomly sampled values are drawn in the range of the uncertainty of each quantity.
@@
With the above input quantities, we track the number of YSOs found in each class as a function of time. Figure 4 (left panel) displays four examples (out of 250 generated models) for the case with a flat SFH, a Milky-like IMF, and in each cluster. Figure 4 indicates that within the temporal fluctuations due to the random sampling, the total number of Class 0 protostars is constant over time. The number of Class I and Class F protostars starts to increase after and Myr, respectively, and both reach a plateau after Myr. The numbers of Class II and Class III PMS stars display the same behavior but at later times (i.e., after between two and several million years). The right panel of Fig. 4 displays the time evolution of the number of YSOs in each class for four cases with a Gaussian-like SFH (with the fiducial parameters). This figure clearly shows the peak in Class 0 protostars at Myr and this peak shifts to later epochs for more evolved protostars. WE note that in these cases, the number of Class II and Class III PMS stars also declines at later times because most of the protostars were formed early at Myr and almost no new protostars were formed after Myr.
@@
Using the number of YSOs found in each class and at each epoch, as well as the mean values of their lifetimes (i.e., those suggested by Evans et al. 2009), we can now apply Eq. 1 and calculate the corresponding SFRs. We also measure the true SFR in each model. For this, we take a time step of yr. This guaranties that the smallest protostellar lifetimes are well resolved. We then simply calculated the true SFR as the total mass of protostars that form within each time step:
@@
where runs over all stars that form between and . Figure 5 (left) displays the time evolution of the SFR for four cases where the SFH is flat. We recall that this is usually the assumption that is made in the observations when deriving the SFR. Figure 6 (left panel) displays the true SFR measured directly from the models (dashed lines) and an SFR estimate using the total number of YSOs (full line). For the latter, a characteristic YSO lifetime of 2.5 Myr was used (Pokhrel et al. 2020). The left panels of Fig. 5 and Fig. 6 show that, while the SFR estimates measured using a single class converge to the real value, they do so with a delay that increases for more evolved YSOs. This implies that for cases where the SFH is constant, measurements of the SFR will only be reliable if only Class 0 and I protostars are considered in young star forming regions and Class II PMS stars in older regions (i.e., ages Myr). Figure 6 (left panel) additionally shows that using the total number of YSOs provides a very poor approximation of the real SFR; it underestimates the true SFR by a factor of between and in young star-forming regions and overestimates the true SFR by a factor of a few to several for older regions. The right panels in Fig. 5 and Fig. 6 display the same SFRs calculated for cases where the SFH is a Gaussian-like function with the fiducial values of the parameters, Myr and Myr. The same patterns that are observed with a constant SFH are also prevalent here, with the additional effect that at advanced epochs (i.e., ages Myr), the estimates of the SFR that are either based on the Class II and Class III PMS stars or on the total number of YSOs overestimate the true SFRs by several orders of magnitude. This is due to the fact that while star formation has ceased, Class II and Class III PMS stars are long lived and their numbers bias the measurement of the SFR.
@@
In order to get a clearer picture of the offset between the SFRs estimated using the YSO populations and the true SFR, we show the ratio between these quantities in Fig. 7. In both panels of this figure, the ratios were obtained by averaging over the 250 realizations (i.e., 250 clusters), each with and a Milky Way-like IMF. The top panel corresponds to cases with a flat SFH and the bottom panel to cases with the Gaussian-like SFH, with all clusters in the latter case having Myr and Myr. The light dashed line in both panels corresponds to the 1 Poisson uncertainty on each measurement. Ideally, for these measurements to be accurate, the offset parameter =SFRi/SFRreal would have to be close to the unity. However, this is far from being the case, with the worst departure from unity corresponding to the time-dependent, Gaussian-like SFH, especially around and after the peak of star formation at Myr. For these time-dependent models of the SFH, we expect the level of departure from unity to depend on the duration of the burst (i.e., the width of the Gaussian). We therefore ran additional models, each with 250 clusters with Gaussian-like SFHs. As our fiducial case has a value of Myr, we considered additional cases where Myr (a narrow, high-amplitude burst) and Myr (an extended, low-amplitude burst). The results for these models are displayed in Fig. 8 for the narrow burst (top panel) and the extended burst (bottom panel). Qualitatively, both cases display the same features and are similar to the fiducial case displayed in Fig. 7 (top panel). However, a noticeable difference for the narrow-burst case (i.e., case with Myr) is that all SFR measurements using the different classes of YSOs or their total number, with the exception of the one based on the Class 0 population, fail to reproduce the true SFR.
@@
The conclusion that can be drawn from Fig. 7 and Fig. 8 is that the only reliable estimate of the SFR at all epochs is the one based on the population of Class 0 protostars. The latter measurement is always accurate, irrespective of the SFH. Measurements of the SFR based on the population of more evolved YSOs can provide reliable estimates of the SFR in cases where the SFH is constant or displays a weak burst, and this is only valid when the star-forming regions are themselves more evolved (i.e., with ages of 2 Myr or older, and harboring a significant fraction of evolved YSOs). In all circumstances, estimates of the SFR that rely on the total number of YSOs and a characteristic YSO lifetime (in this work, Myr) are not accurate and this is true irrespective of the SFH and the considered age of the region. Finally, we should mention that we ran models with lower and higher numbers of YSOs, namely and , respectively, while keeping Myr in all models. Models with lower or higher numbers of protostars display the same effects as those observed in models with YSOs, but with an increased and decreased level of temporal fluctuation, respectively (see Appendix A).
@@ 3.2 The effect of the protostellar mass function
Recent studies suggest a non-negligible degree of variation in the parameters that characterize the shape of the IMF of young clusters in the Milky Way and in M31 (Dib 2014; Weisz et al. 2015; Lim et al. 2015; Dib et al. 2017, Dib 2023b). Dib et al. (2017) found that an intrinsic scatter in the IMF parameters is necessary in order to match the observed fraction of single O stars in young Galactic clusters (cluster ages Myr). For a TPL function representing the system IMF, Dib et al. (2017) found that the standard deviation in , , and is , , and M⊙. Here, as we are dealing with a single-star protostellar mass function, we model the distribution of each of the mass function parameters with a Gaussian centered around the Galactic values (, , and M⊙) and with a standard deviation on each parameter that is of its mean value, namely , , and M⊙. Similar to the previous calculations, we generate a sample of 250 clusters, each with and where the parameters of the IMF of each cluster are randomly drawn from these Gaussian distributions. The IMF parameters are randomly drawn from their Gaussian distribution within the uncertainty range. Figure 9 displays the cases for four clusters with such a variable IMF. We then follow the same procedure as above and derive the SFRs based on the different YSO classes and an SFR based on the total YSO population. Adopting the case of a flat SFH, Fig. 10 displays the time evolution of these SFRs normalized by the true SFR. Comparing this figure with Fig. 7 (top panel), we conclude that a variable IMF has little effect on the derived SFRs. The same features observed in the top panel of Fig. 7 and where the IMF of clusters are randomly drawn from a Milky Way-like IMF are seen in Fig. 10; that is, we see a faithful reproduction of the true SFR (SFRreal) by the SFR based on the Class 0 population, and a reproduction of the SFRreal by the SFRs based on the more evolved populations (Class I and beyond) at more advanced epochs. The SFR derived using the total YSO population remains a poor approximation of the true SFR, as it severely underestimates SFRreal at early epochs and overestimates it at later epochs.
@@ 4 Constraining the SFH using protostellar counts and SFR estimates
In §. 3, we argue that the SFR derived using the Class 0 YSO count provides a reliable estimate of the true SFR in star-forming regions, irrespective of the SFH. However, as exemplified by the models with stars, the number of Class 0 YSOs remains quite modest in most, if not all, epochs. Furthermore, Class 0 YSOs are sometimes difficult to detect as they are deeply embedded in the clouds (Evans et al. 2009). In contrast, the numbers of evolved YSOs is much larger. There are approximately 20 times more Class II YSOs than Class 0 YSOs. For this reason, using Class II objects to derive the SFR can be an appealing option in order to make use of the larger number of Class II objects. A caveat here is that the offset between the Class II-derived SFR (SFRII) and the true SFR depends on the SFH. As shown in §. 3, for the case of a flat SFH, the offset between SFRII and SFRreal exists only at early epochs ( 2 Myr from the start of star formation). For the case of a burst-like SFH, the situation is more complex, and an offset between SFRII and SFRreal exists at all times. It is possible to reverse the arguments and assumes that the observed SFRII corresponds to the SFRreal and verify whether this assumption holds for the predictions of a constant SFH or a burst-like one. Figure 11 displays the Class I and Class II SFR offsets plotted against the ratios and , respectively, for the case with a flat SFH. Figure 12 is similar to Fig. 11 but for the fiducial case with a Gaussian-like SFH. The dependencies between the offsets and and the number ratios are noticeably different between these cases with a different SFH.
@@
We can test the usefulness of these relations in discriminating between SFHs by including observational data. We use the data presented in Tables 3 and 5 of Evans et al. (2009) for five nearby molecular clouds, namely Cha II, Lupus, Perseus, Serpens, and Ophiucus. Table 5 of Evans et al. (2009) lists the numbers of YSOs in Class I, the F Class, Class II, and Class III. As the numbers of Class II YSOs are larger than those in all other classes, this indicates that star formation started at least Myr ago, which is roughly equal to plus a fraction of the Class II lifetime. Table 3 of Evans et al. (2009) also lists SFR estimates for these clouds. These SFR measurements are derived using the Class II YSO populations and were presented in separate works, but are all based on c2d observations. Using the number counts of Class I and Class II YSOs, we use Eq. 1 to calculate SFRI and SFRII and consequently and . If the SFR values quoted for these clouds are the true values, the ratio for these clouds should coincide with the prediction of the model with a constant SFR. The results shown in Fig. 11 (bottom panel) confirm that this is indeed the case. Furthermore, the upper panel of Fig. 11 shows that the versus relation for these five clouds is also reasonably well matched by the model with a constant SFH. In contrast, the model with a burst-like SFH does not overlap with the observational data, as can be seen in the lower and upper panels of Fig. 12. The conclusion that can be drawn from Figs. 11 and 12 is that the SFH in these five nearby and low- to intermediate-mass clouds (from to M⊙; Table 1 in Evans et al. 2009 and references therein) is relatively constant. It would be interesting to include data points for more massive star-forming regions and regions that are located further away in the Galaxy.
@@
and for the time-dependent, Gaussian-like SFH, we find:
As pointed out in §. 2.2, the Class III YSOs observed by Spitzer for these nearby star-forming regions are incomplete and Eqs. 6 and 7 should be revised with more complete class III censuses. More complete censuses of Class III YSOs could be provided by X-ray observations, as shown by Preibisch et al. (2005) in the Orion Nebular Cluster. However, we note that more accurately accounting for Class III YSOs will reinforce our conclusions concerning the comparison of our models with the observations. Accounting for a higher number of missing Class III YSOs would cause the observational number ratios and in Figs. 11 and 12 to move to the left. This would not affect the agreement between the flat SFH models and the observations, but it would cause burst-like models to deviate further from the observations. While it is not yet possible to include a more massive region, such as the entire Orion cloud, because the census of Class II YSOs is incomplete (Furlan et al. (2016)444Megeath et al. (2012,2016) performed a comprehensive survey of the YSO population in the Orion clouds. However the YSOs in these works were categorized in the broad category of protostars and PMS stars and for this reason, the Orion data points could not be included in Figs. 11 and 12. Furlan et al. (2016) found that the number of Class 0 YSOs in Orion (). Using Eq. 1, we conclude that the true SFR for Orion is M⊙ yr-1. This value sits in the middle between the SFRs measured by Lada et al. (2010) for the Orion A and Orion B clouds.